OpenStudy (anonymous):

If cos x = .34, what is sin x? a .94 b .79 c .34 d .57

OpenStudy (phi):

are you allowed a calculator ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

?

OpenStudy (asnaseer):

use the identity:\[\cos^2(x)+\sin^2(x)=1\]to find \(\sin(x)\)